# PolypDB: A Curated Multi-Center Dataset for Development of AI Algorithms in Colonoscopy

### v1

@@ Abstract
Colonoscopy is the primary method for examination, detection, and removal of polyps. Regular screening helps detect and prevent colorectal cancer at an early curable stage. However, challenges such as variation among the endoscopists’ skills, bowel quality preparation, and complex nature of the large intestine which cause large number of polyp miss-rate. These missed polyps can develop into cancer later on, which underscores the importance of improving the detection methods. A computer-aided diagnosis system can support physicians by assisting in detecting overlooked polyps. However, one of the important challenges for developing novel deep learning models for automatic polyp detection and segmentation is the lack of publicly available, multi-center large and diverse datasets. To address this gap, we introduce PolypDB, a large scale publicly available dataset that contains 3934 still polyp images and their corresponding ground truth from real colonoscopy videos to design efficient polyp detection and segmentation architectures. The dataset has been developed and verified by a team of 10 gastroenterologists. PolypDB comprises of images from five modalities: Blue Light Imaging (BLI), Flexible Imaging Color Enhancement (FICE), Linked Color Imaging (LCI), Narrow Band Imaging (NBI), and White Light Imaging (WLI) and three medical centers from Norway, Sweden and Vietnam. Thus, we split the dataset based on modality and medical center for modality-wise and center-wise analysis. We provide a benchmark on each modality using eight popular segmentation methods and six standard benchmark polyp detection methods. Furthermore, we also provide benchmark on center-wise under federated learning settings. Our dataset is public and can be downloaded at https://osf.io/pr7ms/. More information about the dataset, segmentation, detection, federated learning benchmark and train-test split can be found in https://github.com/DebeshJha/PolypDB.
@@ I Introduction
Colorectal cancer (CRC) represent the third highest cancer incidence and is the second most common cause of cancer-related death worldwide. In 2020, approximately 1.9 million new cases of CRC was detected and causing approximately 935,000 deaths [1]. The relative five-year survival rate for persons younger than 64 years is 68.8% [2]. Colonoscopy is the golden standard for detecting CRC and removal of precancerous lesions such as polyps and very early CRCs. However, colonoscopy is an operator-dependent procedure causing an important variation in polyp detection [3]. Smaller polyps, diminutive ($\leq 5mm$) or (6 to 9mm) sized colon polyps are often missed by the endoscopists. The adenoma miss-rate is reported to be 20%–24% [4] and some missed polyps develop into CRC later on called postcolonoscopy CRC or interval cancer [5] . For a couple of years, computer aided detection (CADe) systems for polyp detection are commercially available and have shown to increase the ADR but the polyps are just marked with a bounding box and do not help the endoscopists to delineate the polyp and to confirm complete resection of the polyp, essential to avoid recurrence and potentially post colonoscopy CRC.
TABLE I: An overview of colon polyp datasets with minimum of 1000 samples.
@@
In this perspective an exact delineation of the polyp would be very helpful. Accurate polyp segmentation is a challenging process. This is because (i) polyp changes their characteristics over time during their development stage, and (ii) shape, size, colors, and appearance of polyps may be very similar to the surrounding mucosa. In some polyp cases, there is a strong camouflage that might trick the endoscopists. Even state-of-the-art (SOTA) deep learning algorithms show false positives for such examples. (iii) The imaging device introduces artifacts like blurriness, flares and lightning conditions that also affect colonoscopy procedure, for example, objects too close to the camera, under or over scene lightning, low resolution of capsular endoscopes, overexposure, reflection from the bright spot, low contrast areas. All these can affect colonoscopy procedure and limit accurate polyp segmentation. (iv) presence of surgical instrument and intestinal residue can also affect accurate polyp segmentation [13].
@@
It is important to improve the diagnostic performance of the colonoscopy procedure. Fulfilling the gap between expert and non-expert endoscopists in detecting and diagnosing colon polyps, one of the significant challenges in colonoscopy [14, 15]. An automatic polyp segmentation algorithm can highlight the potential presence of polyps with pixel-level accuracy and could help endoscopists. Most of the above methods perform reasonably well on the large adenomas ($\geq 10mm$), which are easy to segment while overlooking small, diminutive, and even flat large serrated lesions the main reason for right sided pots colonoscopy colorectal cancer [16]. These types of polyps are challenging to detect even for experienced endoscopists [17]. The performance of these methods is only reported on a single-center dataset. Although they might obtain high performance on the particular test set, they suffer from huge performance gaps when tested on out-of-sample datasets, or datasets collected from different cohort populations and hospitals, captured using different types of scopes and modality leading to generalization failure. Training algorithms on multi-center datasets can improve generalizability and robustness of the network.
@@
The main motivation of our work is to develop and publicly release a large-scale multi-center polyp segmentation and detection dataset to develop advanced computer aided diagnosis (CAD) systems that are robust and generalizable for polyp segmentation method useful for integration into the clinical settings. Our dataset consists of a diverse set of images covering the global representativeness of the population and their annotations useful for performance evaluation and comparison of different Deep learning (DL) based algorithms. Our multi-center dataset consists of datasets from a variety of sources, imaging modalities ( Blue laser imaging (BLI), Flexible spectral Imaging Color Enhancement (FICE), white light imaging (WLI), linked color imaging (LCI)), populations (Norway, Vietnam, Sweden), acquisition protocols (Fujinon system, Olympus) and imaging conditions captured by multi-national expert that are better for early polyp diagnosis. Furthermore, we exploit this multi-center dataset and propose to deveop novel techincal for polyp detection and segmentation. The main contribution of this work are as follow:
@@
    PolypDB— We develop and publicly release a multi-center, multi-modality polyp segmentation and detection dataset that consists of 3934 polyp images, pixel-precise ground truth and bounding box annotations collected from medical centers in Norway, Sweden and Vietnam.
@@
    First-ever open access multi-modality dataset — PolypDB consists of five distinct modalities such as BLI, FICE, LCI, NBI and WLI. This is the first-ever open access dataset to feature five distinct modalities along by gastroenterologist-verified ground truth.
@@
    Baseline benchmark — We evaluated PolypDB on each modality using eight segmentation methods, six object detection methods, and six federated learning approaches, establishing a robust baseline benchmark.
@@ II-A Study design
The PolypDB dataset is a collection of colonoscopy examination images from three medical hospitals in Norway, Sweden, and Vietnam. Figure 3 presents the example images from BLI, FICE, LCI, NBI and WLI along with their corresponding bounding box ground truth and color coded segmentation masks. The main motivation behind the development of the dataset is the imperative need for early detection of early diagnosis of CRC precursors, aimed to reduce the incidence. Although some publicly available datasets already exists (please refer to Table I), there is no modality wise dataset till date. Also multi-center open access dataset is limited in the community. Therefore, a new multi-modality and multi-center dataset holds a significant relevance. This diversity is crucial due to regional and demographic disparities in CRC incidence rates. The collection of multi-center data increases the diversity and broadens population representation. Additionally, having a multi-center dataset also allows for the inclusion of different types of equipment and imaging protocols, which can also improve the robustness and generalizability of the CAD system leading to better patient outcomes.
II-B Dataset acquisition: Inclusion and exclusion criteria of the images
@@ II-B1 Inclusion criteria
The inclusion criteria of the polyp frames are as follows:
  • Images with polyp(s) in WLI mode and FICE mode.
  • Images with a minimum resolution of $1280\times 720$ pixels.
  • Polyp’s boundary must be clear and well-defined.
  • Boston Bowel Preparation Score (BBPS) $\geq 2$.
  • Images captured in magnification mode.
  • Images with poor quality.
@@ II-B2 Exclusion criteria
The exclusion criteria for the polyp frames are as follows:
  • Already resected (removed) polyps – dead polyp and resection site are visible.
  • Polyps resected and transported in a net.
  • Non-colon polyps (eg stomach or small bowel).
  • Polyps injected with blue dye and snare around polyp neck.
  • Resection sites covered in blood, where residual polyps are unclear.
  • Unclear if polyp or stool remnants.
  • Normal anatomical structure.
  • Images in magnification mode.
  • Images with poor quality.
  • Blurry, shaky.
  • Too dark or having flare.
  • Have much liquid (faeces, blood) and mucus.
  • Image with already resected polyps or resection sites.
  • Images containing endoscopic tools such as caps, injection needles, snares, biopsy forceps, and clips.
@@ II-C1 Baerum Hospital, Vestre Viken Hospital Trust, Bærum, Norway (Center 1)
The polyp images were collected and verified by experienced gastroenterologists from Vestre Viken Health Trust in Norway. Some images have been collected from the unlabelled class of HyperKvasir dataset [7]. There are 99,417 endoscopic frames in HyperKvasir dataset. We identified 3000 WLI polyps frames, label them and sent it our our gastroenterologists. Out of 3000 images, only 2588 were incorporate them into our datasets. Others were excluded based on the exclusion criteria.
Additionally, we also selected 136 NBI images from Unlabelled HyperKvasir class. We curated the ground truth for both WLI and NBI which was verified by a team of expert gastroenterologists. By labelling such datasets, we are making use of unlabelled frames, which were never explored for the development of new tools.
@@ II-C2 Karolinska University Hospital, Stockholm, Sweden (Center 2)
The images were collected and verified by an experienced gastroenterologist (10+ years of experience) from Karolinska Medical Hospital in Sweden. Although from their center, we received images from entire GI tract, the number of polyp images were relatively limited. Based on exclusion criteria, we selected only 30 WLI polyp images and 10 NBI polyp images from Karolinska. All of these images were completely anonymized according to GDPR requirements for full anonymization.
The dataset BKAI consisted of 1200 endoscopic images with polyps in 4 light modes: WLI, LCI, BLI and FICE. The data acquisition procedure for both centers are identical and they have same cohort population for examination. Therefore, we consider them as the single center in this study given that both centers are located in the same city. Out of total of 1200 images, 600 images were obtained from HMU, while other 600 images were sourced from IGH. Specifically, BKAI consists of 1000 WLI polyp images, 60 LCI, 70 FICE and 70 BLI images. These images were labelled and annotated by three expert endoscopists with an experience more than 10 years in Vietnam. We provide both bounding box information and pixel-precise annotation for all of these images so that the dataset can be useful for both object detection and segmentation tasks. We also organized the dataset in center-wise and modality-wise so that it could be useful to facilitate the research in towards specific objective in multiple direction.
II-D Annotation strategies and quality assurance
A team of 10 gastroenterologists (with most of them over 10 years of experience in colonoscopy) and one experienced senior research associate were involved in the data annotation, sorting and the review process of the quality of annotations. The annotations were performed by a senior research associate who has extensive experience in data curation and development using online annotation tool called Labelbox https://labelbox.com/. All images were uploaded to Labelbox, and each frame was labelled considering the reason of interest (area covered by polyp), and the ground truth for each sample was created. Each annotation was cross-verified by at least two senior gastroenterologists. Furthermore, we assign an independent reviewer (senior gastroenterologist) to review all 3934 images. All of the images were annotated by one researcher using the Wacom Cintiq tablet to minimize the heterogeneity in the manual delineation process.
During the review process, the gastroenterologists marked if the frame represented colon polyps and should be included in the dataset. After that, they checked if the annotations for each polyp in a frame were “correct” and clinically acceptable. Finally, the non-polyp images were removed, and annotations were adjusted for incorrect annotations. For modality-wise organization, we provide “images”, “corresponding groundtruth masks” in the segmentation folder and “images” and “corresponding bounding box information” in the detection folder for each modality. The images and corresponding groundtruth contain the same filename.
For the center-wise data organization, we divide the dataset into three centers: Simula, Karolinska, and BKAI. Each center has images, segmentation ground truth, and bounding box information useful for segmentation and polyp detection tasks. All the images are encoded using JPEG compression. Our datasets can be easily downloaded for research and educational purposes. We also encourage the use of our dataset for industrial applications. However, prior consent is required.
@@ II-D Annotation strategies and quality assurance
Exempted Written informed consent
@@
Hanoi Medical University (HMU), Hanoi Vietnam
@@ II-E Ethical and privacy aspects of the data
PolypDB was collected from 3 different medical hospitals. Each center handled the legal, ethical and privacy aspect of the dataset from their center. The data collection center handled all or two of these steps before providing the dataset from the center. Additionally, we believe releasing these datasets would help in the technological development, for example, the development of robust CAD system for polyps and there is a high potential benefit compared to the potential risk. Therefore, we make this dataset public after carefully considering ethical and privacy issues. Table II illustrates the ethical and legal processes fulfilled by each center, along with the endoscopy equipment and recorders used for the data collection.
  • 1.
    Informed consent from the patient was obtained when required. Approval from the institution was always obtained. This also included the purpose of the study and how their datasets will be used.
  • 2.
    Review and approval of the collected data from data inspector, institutional review board or local medical ethics committee depending on their country’s regulations.
  • 3.
    De-identification of the colonoscopy frame prior to release by following laws and regulations related to data privacy and protection in their nation.
@@ III-A1 Dataset
We have conducted experiments in two different settings: (i) modality-wise and (ii) center-wise. For modality-wise settings, we have 3558 WLI polyp images, 146 NBI images, 60 LCI images, 70 BLI, and 70 FICE images. We only experiment with WLI images for center-wise settings because it is common in all three centers. Although, there are 136 NBI polyp images in center 1 and 10 polyp images in center 2. Due to the minimal number of images present in both centers, we exclude it from the experiment.

### v2

@@ Abstract
Colonoscopy is the primary method for examination, detection, and removal of polyps. However, challenges such as variations among the endoscopists’ skills, bowel quality preparation, and the complex nature of the large intestine contribute to high polyp miss-rate. These missed polyps can develop into cancer later, underscoring the importance of improving the detection methods. To address this gap of lack of publicly available, multi-center large and diverse datasets for developing automatic methods for polyp detection and segmentation, we introduce PolypDB, a large scale publicly available dataset that contains 3934 still polyp images and their corresponding ground truth from real colonoscopy videos. PolypDB comprises images from five modalities: Blue Light Imaging (BLI), Flexible Imaging Color Enhancement (FICE), Linked Color Imaging (LCI), Narrow Band Imaging (NBI), and White Light Imaging (WLI) from three medical centers in Norway, Sweden, and Vietnam. We provide a benchmark on each modality and center, including federated learning settings using popular segmentation and detection benchmarks. PolypDB is public and can be downloaded at https://osf.io/pr7ms/. More information about the dataset, segmentation, detection, federated learning benchmark and train-test split can be found at https://github.com/DebeshJha/PolypDB.
@@ I Introduction
Colorectal cancer (CRC) represents the third highest cancer incidence and is the second most common cause of cancer-related death worldwide. In 2020, approximately 1.9 million new cases of CRC were detected, causing approximately 935,000 deaths [1]. The relative five-year survival rate for persons younger than 64 years is 68.8% [2]. Colonoscopy is the gold standard for detecting CRC and removal of precancerous lesions such as polyps and very early CRCs. However, colonoscopy is an operator-dependent procedure causing a significant variation in polyp detection [3]. Smaller polyps, diminutive ($\leq 5\,\text{mm}$) or (6 to 9 mm) sized colon polyps are often missed by the endoscopists. The adenoma miss-rate is reported to be 20%–24% [4] and some missed polyps develop into CRC later on called postcolonoscopy CRC or interval cancer [5]. For a couple of years, computer-aided detection (CADe) systems for polyp detection are commercially available and have shown to increase the adenoma detection rate but the polyps are just marked with a bounding box and do not help the endoscopists to delineate the polyp and confirm complete resection of the polyp, essential to avoid recurrence and potentially post colonoscopy CRC risks.
TABLE I: An overview of colon polyp datasets with a minimum of 1000 samples.
@@
Precise delineation of polyps may be very helpful, especially for difficult ones, such as sessile serrated lesions (SSL). Accurate polyp segmentation is challenging because (i) polyp changes their characteristics over time during their development stage, (ii) their shape, size, colors, and appearance may be very similar to the surrounding mucosa, (iii) In some cases, there is a mucous covering the polyp acting as camouflage that might trick the endoscopists, even with  state-of-the-art (SOTA) deep learning algorithms showing false positives, (iv) imaging device introducing artifacts such as blurriness, flares, and lightning conditions that also affect the colonoscopy procedure, for example, objects too close to the camera, under or over scene lightning, low resolution of capsular endoscopes, overexposure, reflection from the bright spot, low contrast areas and (v) the presence of surgical instruments and intestinal residue can also affect accurate polyp segmentation [13]. All these can affect colonoscopy procedures and limit accurate polyp segmentation and detection.
@@
Fulfilling the gap between expert and non-expert endoscopists in detecting and diagnosing colon polyps is one of the most critical challenges in colonoscopy [14, 15]. Most of the DL methods perform reasonably well on the large adenomas ($\geq 10\,\text{mm}$), which are easy to segment while overlooking small, diminutive, and even flat large SSLs, the main reason for right-sided post colonoscopy colorectal cancer [16]. However, SSLs are challenging to detect and delineate even for experienced endoscopists [17]. Training DL algorithms on multi-center datasets can improve the generalizability and robustness of the network.
@@
The main motivation of our work is to develop and publicly release a large-scale, multi-center polyp segmentation and detection dataset to be developed to support computer aided diagnosis (CAD) systems that are robust and generalizable for polyp segmentation and detection methods useful for integration into clinical settings. PolypDB consists of a diverse set of annotated images covering the global representativeness of the population and their annotations useful for performance evaluation and comparison of different Deep learning (DL) based algorithms. Our multi-center dataset consists of data from a variety of sources, imaging modalities ( Blue laser imaging (BLI), Flexible spectral Imaging Color Enhancement (FICE), white light imaging (WLI), linked color imaging (LCI)), populations (Norway, Vietnam, Sweden), acquisition protocols (Fujinon system, Olympus) and imaging conditions captured by a multi-national expert that are better for early polyp diagnosis. Furthermore, we exploit this multi-center dataset and propose developing new benchmarks for polyp detection and segmentation both modality and center-wise. The main contributions of this work are as follows:
@@
    PolypDB — We present PolypDB, a multi-center, multi-modality polyp segmentation and detection dataset that consists of 3934 polyp images, pixel-precise ground truth and bounding box annotations collected from medical centers in Norway, Sweden and Vietnam. The diverse dataset helps the model enhance training and testing under real-world conditions.
@@
    First-ever open access multi-modality dataset — PolypDB consists of five distinct modalities such as BLI, FICE, LCI, NBI and WLI. This is the first-ever open-access dataset to feature five distinct modalities along with gastroenterologist-verified ground truth.
@@
    Baseline benchmark — We evaluated PolypDB on each modality using eight segmentation methods, five object detection methods, and six federated learning approaches, establishing a robust baseline benchmark.
@@ II-A Study design
PolypDB is a collection of colonoscopy examination images from three medical hospitals in Norway, Sweden, and Vietnam. Figure 1 presents the example images from different brands Fujifilm, Olympus, Pentax, primarily WLI, and with different types of digital staining such as BLI, FICE, LCI, and NBI, along with their corresponding bounding box ground truth and color coded segmentation masks. PolypDB is developed to address the urgent need for early detection and diagnosis of CRC precursors to reduce the incidence of CRC. Although some publicly available datasets exist (Table I), there is no comprehensive modality-specific dataset to date. Also, multi-center open-access dataset is limited in the community. The multi-modality and multi-center dataset captures the regional and demographic disparities in CRC incidence rates, enhancing data diversity and broadening population representation. Additionally, having a multi-center dataset allows for the inclusion of different types of equipment and imaging protocols, which can also improve the robustness and generalizability of the CAD system, leading to better patient outcomes.
II-B Dataset acquisition: Inclusion and exclusion criteria of the colonoscopy frames
@@ II-B1 Inclusion criteria
The inclusion criteria for the polyp frames are as follows: Images with native colorectal polyp(s) in WLI mode and FICE mode, having a minimum resolution of $1280\times 720$ pixels, and polyp’s boundary must be clear and well-defined. Additionally, Boston Bowel Preparation Score (BBPS) $\geq 2$ and image should be captured in magnification mode.
@@ II-B2 Exclusion criteria
The exclusion criteria ensure that we have high-quality and clinically relevant frames. We excluded frames where a polyp was resected (removed) or transported in a net, those with poor image quality, polyps injected with blue dye and snare around the polyp neck, and resection sites covered in blood, where residual polyps are unclear. Additionally, images were removed if it was unclear if a polyp or stool remnants, those showing normal anatomical structure or images in other image-enhanced modes (BLI, LCI), images in magnification mode. Images with poor quality, such as blurry, shaky, too dark, having a flare or having much liquid (feces, blood) and mucus, were removed. Images with already resected polyps or resection sites, images of polyps with submucosal injection, and images containing endoscopic tools such as caps, injection needles, snares, biopsy forceps, and clips were also excluded.
@@ II-C1 Baerum Hospital, Vestre Viken Hospital Trust, Bærum, Norway (Center 1)
The polyp images were collected and verified by experienced gastroenterologists from Bærum Hospital, Vestre Viken Hospital Trust, Bærum, Norway hospital trust in Norway. Some images have been collected from the unlabelled class of HyperKvasir dataset [7]. There are 99,417 endoscopic frames in HyperKvasir dataset. We identified 3000 WLI polyps frames, labelled them and sent them to our gastroenterologists. Out of 3000 images, only 2588 were incorporated into our datasets. Others were excluded based on the exclusion criteria. Additionally, we selected 136 NBI images from the unlabelled HyperKvasir class. We curated the ground truth for both WLI and NBI, which was verified by a team of expert gastroenterologists. By labelling such datasets, we are making use of unlabelled frames, which were never explored for the development of new tools.
@@ II-C2 Karolinska University Hospital, Stockholm, Sweden (Center 2)
The images were collected and verified by an experienced gastroenterologist (10+ years of experience) from Karolinska University Hospital, Stockholm, Sweden Medical Hospital in Sweden. Although from their center, we received images from the entire GI tract, the number of polyp images was relatively limited. Based on exclusion criteria, we selected only 30 WLI polyp images and 10 NBI polyp images from Karolinska University Hospital, Stockholm, Sweden hospital. All these images were completely anonymized according to GDPR requirements for full anonymization.
@@
Exempted Not required
@@
Hanoi Medical University (HMU), Hanoi, Vietnam
@@
The dataset consisted of 1200 endoscopic images with polyps in 4 light modes: WLI, LCI, BLI and FICE. The data acquisition procedures for both centers are identical, and they examine similar populations. Therefore, we consider a single center in this study, given that both centers are located in the same city. Out of a total of 1200 images, 600 images were obtained from Hanoi Medical University (HMU), Hanoi, Vietnam, while the other 600 images were sourced from Institute of Gastroenterology and Hepatology (IGH), Hanoi, Vietnam. Specifically, BKAI consists of 1000 WLI polyp images, 60 LCI, 70 FICE and 70 BLI images. These images were labelled and annotated by three expert endoscopists with more than 10 years of experience. We provide both bounding box information and pixel-precise annotation for all images to make the dataset useful for both object detection and segmentation tasks. We also organized the dataset center-wise and modality-wise so that it could be useful to facilitate the research towards specific objectives in multiple directions.
II-D Annotation strategies and quality assurance
A team of 8 gastroenterologists (with most of them over 10 years of experience in colonoscopy) and one experienced senior researcher with a computer science background were involved in the data annotation, sorting, and the review process of the quality of annotations. The annotations were performed by a senior research associate who has extensive experience in data curation and development using online annotation tool called Labelbox https://labelbox.com/. All images were uploaded to Labelbox, and each frame was labelled considering the region of interest (area covered by polyp), and the ground truth for each sample was created. Each annotation was cross-verified by at least two senior gastroenterologists. Furthermore, we assign an independent reviewer (senior gastroenterologist) to review all 3934 images. All of the images were annotated by one researcher using the Wacom Cintiq tablet to minimize the heterogeneity in the manual delineation process.
During the review, the gastroenterologists marked if the frame represented colon polyps and should be included in the dataset. After that, they checked if the annotations for each polyp in a frame were “correct” and clinically acceptable. Finally, the non-polyp images were removed, and annotations were adjusted for incorrect annotations. For modality-wise organization, we provide “images”, “corresponding ground truth masks” in the segmentation folder and “images” and “corresponding bounding box information” in the detection folder for each modality. The images and corresponding ground truth contain the same filename. For the center-wise data organization, we divide the dataset into three centers: Simula, Karolinska, and BKAI. Each center has images, segmentation ground truth, and bounding box information useful for segmentation and polyp detection tasks. All images are encoded using JPEG compression.
@@ II-E Ethical and privacy aspects of the data
The three medical hospitals involved in the PolypDB acquisition handled either all or at least two of the given steps, focusing on legal, ethical and privacy aspects of the dataset. Additionally, we believe releasing these datasets would help in the technological development, for example, the development of robust CAD system for polyps and there is a high potential benefit compared to the potential risk. Therefore, we make this dataset public after carefully considering ethical and privacy issues. Table II illustrates the ethical and legal processes fulfilled by each center, along with the endoscopy equipment and recorders used for the data collection. 1) Informed consent from the patient was obtained when required. Approval from the institution was always obtained. This also included the purpose of the study and how their datasets will be used. 2) Review and approval of the collected data from data inspectorate, institutional review board or local medical ethics committee depending on their country’s regulations. 3) De-identification of the colonoscopy frame prior to the export from the hospitals’ medical records release by following laws and regulations related to data privacy and protection in their nation.
@@ III-A1 Dataset
We have conducted experiments in two different settings: (i) modality-wise and (ii) center-wise. For modality-wise settings, we have 3558 WLI polyp images, 146 NBI images, 60 LCI images, 70 BLI, and 70 FICE images. We only experiment with WLI images for center-wise settings because it is common in all three centers. Although there are 136 NBI polyp images in center 1 and 10 polyp images in center 2, due to the minimal number of images present in both centers, we exclude them from the experiment.